# Nonlinear moving horizon estimation for robust state and parameter estimation – extended version

### v2

@@ 1 Introduction
Contribution: We propose an MHE scheme for joint state and parameter estimation for general nonlinear discrete-time systems subject to process disturbances and measurement noise. Our arguments are based on recent MHE results [2, 13, 26], where only state (but no parameter) estimation is considered. The MHE scheme avoids a uniform PE condition and instead uses online information about the current excitation of the parameters to suitably adjusts the regularization term in the cost function (Section 3). We establish a bound on the state and parameter estimation error that is valid for all times (even if the parameters are never or only rarely excited), and which improves the more often sufficient excitation is present. The bound specializes to a robust global exponential stability property under an additional uniform condition on the maximum duration of insufficient excitation. A method to online monitor the level of excitation that is applicable to general nonlinear systems is provided in Section 4. In combination, we provide a flexible MHE scheme for robust joint state and parameter estimation with intuitive error bounds that are valid independent of the parameter excitation. The numerical example in Section 5 illustrates that the proposed approach is able to efficiently compensate for phases of weak excitation and ensure reliable estimation results for all times.
@@
Notation: The set of integers is denoted by $\mathbb{I}$, the set of all integers greater than or equal to $a$ for any $a\in\mathbb{I}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for any $a,b\in\mathbb{I}$ by $\mathbb{I}_{[a,b]}$. The floor-function applied to some $b\in\mathbb{R}_{\geq 0}$ is defined as $\lfloor b\rfloor=\max\{b^{\prime}\in\mathbb{I}_{\geq 0}:b^{\prime}\leq b\}$. The $n\times n$ identity matrix is denoted by $I_{n}$ and the $n\times m$ zero matrix by $0_{n\times m}$, where we omit the indices if the dimension is unambiguous from the context. The weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$ with respect to a positive definite matrix $Q=Q^{\top}$ is defined as $\|x\|_{Q}=\sqrt{x^{\top}Qx}$ with $\|x\|=\|x\|_{I_{n}}$; the minimal and maximal eigenvalues of $Q$ are denoted by $\underline{\lambda}(Q)$ and $\overline{\lambda}(Q)$, respectively. For two matrices $A=A^{\top}$ and $B=B^{\top}$, we write $A\succeq B$ ($A\succ B$) if $A-B$ is positive semi-definite (positive definite). For $A,B$ positive definite, the maximum generalized eigenvalue (i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$) is denoted by $\overline{\lambda}(A,B)$.
@@ 2 Problem setup and preliminaries
:=$\mathbb{Z}:=\{(x,u,w,p)\in\mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{P}:f(x,u,w,p)\in\mathbb{X}\}$
@@
$U(f(x,u,w,p),f(\tilde{x},u,\tilde{w},\tilde{p}))$
@@
For some fixed matrices $S_{\mathrm{p}},P_{\mathrm{p}},Q_{\mathrm{p}},R_{\mathrm{p}}\succ 0$ and a constant $\eta_{\mathrm{p}}\in[0,1)$, we define
@@
$\in\mathbb{Z}^{T}\times\mathbb{Z}^{T}:$
$x_{t+1}=f(x_{t},u_{t},w_{t},p),\ \tilde{x}_{t+1}=f(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p}),$
$y_{t}=h(x_{t},u_{t},w_{t},p),\ \tilde{y}_{t}=h(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p}),$
@@
The set $\mathbb{E}_{T}$ contains all trajectory pairs that exhibit a $T$-step distinguishability property with respect to the parameters. Specifically, for two trajectories that share the same initial state and the same disturbance inputs, if they form a pair contained in the set $\mathbb{E}_{T}$ it holds that the sum of their output differences is zero if and only if their parameters are the same. In Section 3.3, we discuss the relation between detectability of the states (Assumption 1), excited trajectory pairs (Definition 2), uniform excitation, and a uniform joint detectability condition for both the states and the parameters.
@@ 3.1 Design
Assumption 3 ensures large tuning capabilities of the cost function (5) while satisfying certain relations to the detectability and excitation properties from Assumption 1 and Definition 2. This is conceptually similar to the recent MHE literature (for state estimation) and typically permits a less conservative stability analysis due to the structural similarities between the detectability condition, MHE, and stability, cf., e.g., [26, Sec. III.D] and compare also [10, 13]. Potential time dependency of the weighting matrices in (9)–(12) can be used to incorporate additional knowledge (e.g., by choosing Kalman filter covariance update laws [22, 23]), which can be beneficial in practice to increase estimation performance. Assumption 3 implies that the cost function (5) is radially unbounded in the decision variables, which together with continuity of $f$ and $h$ ensures that the estimation problem (4) and (5) admits a globally optimal solution at any time $t\in\mathbb{I}_{\geq 0}$, cf. [24, Sec. 4.2]. We denote a minimizer by $(\hat{x}^{*}_{t-N_{t}|t},\hat{p}^{*}_{|t},\hat{w}_{\cdot|t}^{*})$, and the corresponding optimal state sequence by $\{\hat{x}^{*}_{j|t}\}_{j=t-N_{t}}^{t}$. The resulting state and parameter estimates at time $t\in\mathbb{I}_{\geq 0}$ are then given by $\hat{x}_{t}=\hat{x}^{*}_{t|t}$ and $\hat{p}_{t}=\hat{p}^{*}_{|t}$, yielding the estimation error
@@ 3.2 Stability Analysis
where $k\in\mathbb{I}_{\geq 0}$, $i_{m}\in\mathbb{I}_{[1,k]}$ for $k\in\mathbb{I}_{\geq 1}$, and $j\in\mathbb{I}_{\geq 0}$ are defined as follows. To this end, let the set $\mathcal{T}_{t}:=\left\{\tau\in\mathbb{I}_{[N,t]}:t-\tau-N\left\lfloor\frac{t-\tau}{N}\right\rfloor=0,X_{\tau}\in\mathbb{E}_{N}\right\}$ contain all past time instances from the set $\{t,t-N,t-2N,...\}$ at which the excitation condition used in (14) was met (in the following also referred to as PE horizons for simplicity). The variable $k$ denotes the total number of PE horizons that occurred until the current time $t$ and is defined as the cardinality of $\mathcal{T}_{t}$, i.e., $k:=|\mathcal{T}_{t}|$. Suppose that $k\in\mathbb{I}_{\geq 1}$. The sequence $\{t_{m}\}_{m=1}^{k}$ contains time instances corresponding to PE horizons, where $t_{1}:=\max\{\tau\in\mathcal{T}_{t}\}$ and $t_{m+1}=\max\{\tau\in\mathcal{T}_{t}:\tau<t_{m}\}$ for $m\in\mathbb{I}_{[1,k-1]}$ if $k\in\mathbb{I}_{\geq 2}$. The sequence $\{i_{m}\}_{m=1}^{k}$ denotes the numbers of non-PE horizons (i.e., past time instances where the excitation condition (14) was not met) between two successive times $t_{m}$ and $t_{m+1}$ with $i_{m}=({t_{m}-N-t_{m+1}})/{N}$ for $m\in\mathbb{I}_{[1,k-1]}$ if $k\in\mathbb{I}_{\geq 2}$, and111 Note that $\{t_{m}\}_{m=1}^{k},\{i_{m}\}_{m=1}^{k}$ do not have to be formally defined for $k=0$, as this yields an empty sum in (27). $i_{k}=(t_{k}-N-l)/N$. Finally, $j$ stands for the number of non-PE horizons that occurred between time $t$ and $t_{1}$ if $k\geq 1$ (and between time $t$ and $l$ if $k=0$), i.e., $j:=({t-\max\{\tau\in\mathcal{T}_{t},l\}})/{N}$. Overall, the partitioning in (27) allows us to theoretically cover occasional occurrence of PE horizons, which of course includes the special cases in which PE horizons never occur ($k=0$), or in which all horizons are PE ($j=0$, $i_{m}=0$ for all $m=1,...,k$).
@@
Let Assumption 1 hold. Consider the MHE scheme (4) with the cost function (5) satisfying Assumption 3. Suppose that the horizon length $N$ satisfies (24)–(26). Then, it holds that
@@
    Theorem 6 provides a bound on the state and parameter estimation error222 By the definition of $\Gamma_{1}$ in (15) and the lower bound in (2). that is valid regardless of the parameter excitation; it also applies if the excitation condition in (14) is never met (which corresponds $k=0$ in (28)) and constitutes a bounded-disturbance bounded-estimation-error property.
@@
Proof of Theorem 6. We prove the statement in five parts—namely, by deriving bounds 1) on $\Gamma_{1}(t,\cdot)$ involving data from $[t-jN,t-1]$; 2) on $\Gamma_{2}(\cdot)$ involving data from $[t_{2},t_{1}-1]$; 3) on $\Gamma_{2}(\cdot)$ involving data from $[t_{k},t_{1}-1]$; 4) on $\Gamma_{2}(\cdot)$ involving data from $[0,l-1]$; 5) on $\Gamma_{1}(t,\cdot)$ involving all available data from $[0,t-1]$.
Part 1. We define $\tilde{\Gamma}_{1}(t):=\Gamma_{1}(t,\hat{x}_{t},x_{t},\hat{p}_{t},p)$ for notational brevity and assume that $j\in\mathbb{I}_{\geq 1}$. From Lemma 4, the fact that $\|\bar{x}_{t-N}-x_{t-N}\|_{\overline{W}}^{2}\leq\overline{\lambda}(\overline{W},\underline{U})U(\bar{x}_{t-N},x_{t-N})$ by (2), and the contraction condition (26), we obtain
@@
where $c_{1}^{\prime}=2c_{1}(N)$, $2c_{1}(N)\eta_{\mathrm{1}}^{-N}\overline{Q}\preceq Q$ with $Q$ from (32), and $\rho$ satisfies $\rho<1$. In (33), note that $\bar{x}_{t-N}=\hat{x}_{t-N}$ and $\bar{p}_{t-N}=\bar{p}_{t-qN}=\bar{p}_{t-jN}$ for all $q=1,...,j$ due to the update rule (14). Since $U(\hat{x}_{t-N},x_{t-N})\leq\Gamma_{1}(t-N,\hat{x}_{t-N},x_{t-N},\hat{p}_{t-N},p)$ by (15), we can recursively apply (33) for $j$ times, yielding
@@
$\sum_{m=1}^{k}\mu^{(m-1)N}\sum_{r={1}}^{(i_{m}+1)N}\bar{\mu}^{r-1}\|w_{t-jN-\sum_{q=1}^{m-1}(i_{q}+1)N-r}\|_{Q}^{2}$
$\leq\sum_{r={1}}^{t-jN-l}\bar{\mu}^{r-1}\|w_{t-jN-r}\|_{Q}^{2}$ (47)
by (45) with $s=m-1$ and (27). Hence, from Theorem 6, the definition of $\mu_{\kappa}$, and (45)–(47), we obtain
@@ 3.3 Special case: uniform persistent excitation
($\Big{(}\{(x_{t},u_{t},w_{t},p)\}_{t=0}^{K-1},\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\}_{t=0}^{K-1}\Big{)}\in\mathbb{E}_{K}$
@@
Proof. Consider some $K\in\mathbb{I}_{\geq 1}$. Let the two sequences $\{(x_{t},u_{t},w_{t},p)\}_{t=0}^{K-1}\in\mathbb{Z}^{K}$ and $\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\}_{t=0}^{K-1}\in\mathbb{Z}^{K}$ satisfy (1) for all $t\in\mathbb{I}_{[0,K-1]}$. Define the corresponding outputs $y_{t}=h(x_{t},u_{t},w_{t},p)$ and $\tilde{y}_{t}=h(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})$, $t\in\mathbb{I}_{[0,K-1]}$. For the sake of conciseness, define $\Delta x_{t}:=x_{t}-\tilde{x}_{t}$ for $t\in\mathbb{I}_{[0,K]}$, $\Delta w_{t}:=w_{t}-\tilde{w}_{t}$ and $\Delta y_{t}:=y_{t}-\tilde{y}_{t}$ for $t\in\mathbb{I}_{[0,K-1]}$, and $\Delta p:=p-\tilde{p}$.
@@
which is an (exponential) i-IOSS bound for the system (1) considering $p$ as an additional constant input. Existence of an i-IOSS Lyapunov function $U(x,\tilde{x})$ and matrices $\underline{U},\overline{U},Q_{\mathrm{x}},R_{\mathrm{x}}\succ 0$ and $\eta_{\mathrm{x}}\in[0,1)$ satisfying Assumption 1 follows by a straightforward extension of the converse Lyapunov theorem from [3]. Now fix some $T\in\mathbb{I}_{\geq 1}$ and consider some $K\in\mathbb{I}_{\geq T}$. From (56) with $t=K$ and the facts that $\eta^{K}\leq\eta^{T}$ and $\|\Delta x_{K}\|^{2}\geq 0$, we obtain
@@
for all $K\in\mathbb{I}_{\geq T}$. Since there always exists $T\in\mathbb{I}_{\geq 1}$ such that $(\underline{\lambda}(\underline{G})-\overline{\lambda}(\overline{G})\eta^{T})>0$, Assumption 9 is satisfied, which finishes this proof. ∎ Proposition 10 essentially implies that state detectability (Assumptions 1) and uniform PE of the parameters (Assumption 9) is equivalent to uniform detectability (exponential i-IOSS) of the augmented state $x_{\mathrm{a}}^{\top}=[x^{\top},p^{\top}]$. Consequently, under these assumptions one could simply consider the augmented state $x_{\mathrm{a}}$ and apply MHE schemes and theory for state estimation (e.g., [2, 13, 26]). However, Assumption 9 is restrictive, usually not satisfied in practice, and its a priori verification is generally impossible. The proposed method from Section 3.1, on the other hand, provides a strict relaxation, since it is applicable in the practically relevant case where the parameters are only rarely (or never) excited (which violates Assumption 9 and hence implies that the augmented state cannot be uniformly detectable (i.e., exponentially i-IOSS) and no joint i-IOSS Lyapunov function satisfying (50) and (51) can exist).

### v3

@@ 1 Introduction
Contribution: We propose an MHE scheme for joint state and parameter estimation for general nonlinear discrete-time systems subject to process disturbances and measurement noise. Our arguments are based on recent MHE results [2, 13, 27], where only state (but no parameter) estimation is considered. The MHE scheme avoids a uniform PE condition and instead uses online information about the current excitation of the parameters to suitably adjust the regularization term in the cost function (Section 3). We establish a bound on the state and parameter estimation error that is valid for all times (even if the parameters are never or only rarely excited), which improves the more often sufficient excitation is present. The bound specializes to a robust global exponential stability property under an additional uniform condition on the maximum duration of insufficient excitation. A method to online monitor the level of excitation that is applicable to general nonlinear systems is provided in Section 4, compare also [26] for further details. In combination, we provide a flexible MHE scheme for robust joint state and parameter estimation with intuitive error bounds that are valid independent of the parameter excitation. The numerical example in Section 5 illustrates that the proposed approach is able to efficiently compensate for phases of weak excitation and ensure reliable estimation results for all times.
@@
Notation: The set of integers is denoted by $\mathbb{I}$, the set of all integers greater than or equal to $a$ for any $a\in\mathbb{I}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for any $a,b\in\mathbb{I}$ by $\mathbb{I}_{[a,b]}$. The floor-function applied to some $b\in\mathbb{R}_{\geq 0}$ is defined as $\lfloor b\rfloor=\max\{b^{\prime}\in\mathbb{I}_{\geq 0}:b^{\prime}\leq b\}$. The $n\times n$ identity matrix is denoted by $I_{n}$ and the $n\times m$ zero matrix by $0_{n\times m}$, where we omit the indices if the dimension is unambiguous from the context. The weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$ with respect to a positive definite matrix $Q=Q^{\top}$ is defined as $\|x\|_{Q}=\sqrt{x^{\top}Qx}$ with $\|x\|=\|x\|_{I_{n}}$; the minimal and maximal eigenvalues of $Q$ are denoted by $\underline{\lambda}(Q)$ and $\overline{\lambda}(Q)$, respectively. For two matrices $A=A^{\top}$ and $B=B^{\top}$, we write $A\succeq B$ ($A\succ B$) if ABA-B is positive semi-definite (positive definite). For $A,B$ positive definite, the maximum generalized eigenvalue (i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$) is denoted by $\overline{\lambda}(A,B)$.
@@ 2 Problem setup and preliminaries
$\mathbb{Z}:=\{(x,u,w,p)\in\mathbb{X}\times\mathbb{U}\times\mathbb{W}\times\mathbb{P}:f(x,u,w,p)\in\mathbb{X}\}$
@@
$)\displaystyle U(f(x,u,w,p),f(\tilde{x},u,\tilde{w},\tilde{p}$))
@@
Fix $S_{\mathrm{p}},P_{\mathrm{p}},Q_{\mathrm{p}},R_{\mathrm{p}}\succ 0$ and $\eta_{\mathrm{p}}\in[0,1)$. We define the set containing all excited trajectory pairs of length $T\in\mathbb{I}_{\geq 0}$ as
$\mathbb{E}_{T}:=\Big\{$ $\left(\left\{(x_{t},u_{t},w_{t},p)\right\}_{t=0}^{T-1},\left\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\right\}_{t=0}^{T-1}\right)$
$\in\mathbb{Z}^{T}\times\mathbb{Z}^{T}:$
@@
[…]
@@
The set $\mathbb{E}_{T}$ contains all trajectory pairs of length $T$ that exhibit a $T$-step distinguishability property with respect to the parameters. Specifically, for two trajectories that share the same initial state and the same disturbance inputs, if they form a pair contained in the set $\mathbb{E}_{T}$, it holds that the sum of their output differences is zero if and only if their parameters are the same. In Section 3.3, we discuss the relation between detectability of the states (Assumption 1), excited trajectory pairs (Definition 2), uniform excitation, and a uniform joint detectability condition for both the states and the parameters.
@@ 3.1 Design
Assumption 3 ensures large tuning capabilities of the cost function (5) while satisfying certain relations to the detectability and excitation properties from Assumption 1 and Definition 2. This is conceptually similar to the recent MHE literature (for state estimation) and typically permits a less conservative stability analysis due to the structural similarities between the detectability condition, MHE, and stability, cf., e.g., [27, Sec. III.D] and compare also [10, 13]. Potential time dependency of the weighting matrices in (9)–(12) can be used to incorporate additional knowledge (e.g., by choosing Kalman filter covariance update laws [22, 23]), which can be beneficial in practice to increase estimation performance. Assumption 3 implies that the cost function (5) is radially unbounded in the decision variables, which together with continuity of $f$ and $h$ ensures that the estimation problem (4) and (5) admits a globally optimal solution at any time $t\in\mathbb{I}_{\geq 0}$, cf. [24, Sec. 4.2]. We denote a minimizer by $(\hat{x}^{*}_{t-N_{t}|t},\hat{p}^{*}_{|t},\hat{w}_{\cdot|t}^{*})$, and the corresponding optimal state sequence by $\{\hat{x}^{*}_{j|t}\}_{j=t-N_{t}}^{t}$. The resulting state and parameter estimates at time $t\in\mathbb{I}_{\geq 0}$ are then given by $\hat{x}_{t}=\hat{x}^{*}_{t|t}$ and $\hat{p}_{t}=\hat{p}^{*}_{|t}$, yielding the estimation error
@@ 3.2 Stability Analysis
where $k\in\mathbb{I}_{\geq 0}$, $i_{m}\in\mathbb{I}_{[1,k]}$ for $k\in\mathbb{I}_{\geq 1}$, and $j\in\mathbb{I}_{\geq 0}$ are defined as follows. To this end, let the set $\mathcal{T}_{t}:=\left\{\tau\in\mathbb{I}_{[N,t]}:t-\tau-N\left\lfloor\frac{t-\tau}{N}\right\rfloor=0,X_{\tau}\in\mathbb{E}_{N}\right\}$ contain all past time instances from the set $\{t,t-N,t-2N,$...\} at which the excitation condition used in (14) was met (in the following also referred to as PE horizons for simplicity). The variable $k$ denotes the total number of PE horizons that occurred until the current time $t$ and is defined as the cardinality of $\mathcal{T}_{t}$, i.e., $k:=|\mathcal{T}_{t}|$. Suppose that $k\in\mathbb{I}_{\geq 1}$. The sequence $\{t_{m}\}_{m=1}^{k}$ contains time instances corresponding to PE horizons, where $t_{1}:=\max\{\tau\in\mathcal{T}_{t}\}$ and $t_{m+1}=\max\{\tau\in\mathcal{T}_{t}:\tau<t_{m}\}$ for $m\in\mathbb{I}_{[1,k-1]}$ if $k\in\mathbb{I}_{\geq 2}$. The sequence $\{i_{m}\}_{m=1}^{k}$ denotes the numbers of non-PE horizons (i.e., past time instances where the excitation condition (14) was not met) between two successive times $t_{m}$ and $t_{m+1}$ with $i_{m}=({t_{m}-N-t_{m+1}})/{N}$ for $m\in\mathbb{I}_{[1,k-1]}$ if $k\in\mathbb{I}_{\geq 2}$, and [^1] Note that $\{t_{m}\}_{m=1}^{k},\{i_{m}\}_{m=1}^{k}$ do not have to be formally defined for $k=0$, as this yields an empty sum in (27). $i_{k}=(t_{k}-N-l)/N$. Finally, $j$ stands for the number of non-PE horizons that occurred between time $t$ and $t_{1}$ if $k\geq 1$ (and between time $t$ and $l$ if $k=0$), i.e., $j:=({t-\max\{\tau\in\mathcal{T}_{t},l\}})/{N}$. Overall, the partitioning in (27) allows us to theoretically cover occasional occurrence of PE horizons, which of course includes the special cases in which PE horizons never occur ($k=0$), or in which all horizons are PE ($j=0$, $i_{m}=0$ for all m=1,,km=1,...,k).
@@
Let Assumption 1 hold. Consider the MHE scheme (4) with the cost function (5) satisfying Assumption 3. Suppose that the horizon length NN satisfies (24)–(26). Then, it holds that
@@
    Theorem 6 provides a bound on the state and parameter estimation error [^2] By the definition of $\Gamma_{1}$ in (15) and the lower bound in (2). that is valid regardless of the parameter excitation; it also applies if the excitation condition in (14) is never met (which corresponds $k=0$ in (28)) and constitutes a bounded-disturbance bounded-estimation-error property.
@@
Proof of Theorem 6. We prove the statement in five parts—namely, by deriving bounds 1) on $\Gamma_{1}(t,\cdot)$ involving data from [tjN,t1][t-jN,t-1]; 2) on $\Gamma_{2}(\cdot)$ involving data from $t_{2},t_{1}-1]$; 3) on $\Gamma_{2}(\cdot)$ involving data from $t_{k},t_{1}-1]$; 4) on $\Gamma_{2}(\cdot)$ involving data from [0,l1][0,l-1]; 5) on $\Gamma_{1}(t,\cdot)$ involving all available data from [0,t1][0,t-1].
Part 1. We define $\tilde{\Gamma}_{1}(t):=\Gamma_{1}(t,\hat{x}_{t},x_{t},\hat{p}_{t},p)$ for notational brevity and assume that $j\in\mathbb{I}_{\geq 1}$. By applying Lemma 4 together with the fact that $\|\bar{x}_{t-N}-x_{t-N}\|_{\overline{W}}^{2}\leq\overline{\lambda}(\overline{W},\underline{U})U(\bar{x}_{t-N},x_{t-N})$ by (2) and the contraction condition (26), we obtain
@@
where $c_{1}^{\prime}=2c_{1}(N)$, $2c_{1}(N)\eta_{\mathrm{1}}^{-N}\overline{Q}\preceq Q$ with $Q$ from (32), and $\rho$ satisfies $\rho<1$. In (33), note that $\bar{x}_{t-N}=\hat{x}_{t-N}$ and $\bar{p}_{t-N}=\bar{p}_{t-qN}=\bar{p}_{t-jN}$ for all q=1,,jq=1,...,j due to the update rule (14). Since $U(\hat{x}_{t-N},x_{t-N})\leq\Gamma_{1}(t-N,\hat{x}_{t-N},x_{t-N},\hat{p}_{t-N},p)$ by (15), we can recursively apply (33) for $j$ times, yielding
@@
$\sum_{m=1}^{k}\hskip-2.15277pt\mu^{(m-1)N}\hskip-4.30554pt\sum_{r={1}}^{(i_{m}+1)N}\hskip-6.45831pt\bar{\mu}^{r-1}\|w_{t-jN-\sum_{q=1}^{m-1}(i_{q}+1)N-r}\|_{Q}^{2}$
$\leq\sum_{r={1}}^{t-jN-l}\hskip-6.45831pt\bar{\mu}^{r-1}\|w_{t-jN-r}\|_{Q}^{2}$ (47)
by (45) with s=m1s=m-1 and (27). Hence, from Theorem 6, the definition of $\mu_{\kappa}$, and (45)–(47), we obtain
@@ 3.3 Special case: uniform persistent excitation
$\Big(\{(x_{t},u_{t},w_{t},p)\}_{t=0}^{K-1},\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\}_{t=0}^{K-1}\Big)\in\mathbb{E}_{K}$
@@
[…]
@@
which is an (exponential) i-IOSS bound for the system (1) considering $p$ as an additional constant input. Existence of an i-IOSS Lyapunov function U(x,x~)U($\tilde{x}$) and matrices $\underline{U},\overline{U},Q_{\mathrm{x}},R_{\mathrm{x}}\succ 0$ and $\eta_{\mathrm{x}}\in[0,1)$ satisfying Assumption 1 follows by a straightforward extension of the converse Lyapunov theorem from [3]. Now fix some $T\in\mathbb{I}_{\geq 1}$ and consider some $K\in\mathbb{I}_{\geq T}$. From (56) with $t=K$ and the facts that $\eta^{K}\leq\eta^{T}$ and $\|\Delta x_{K}\|^{2}\geq 0$, we obtain
@@
for all $K\in\mathbb{I}_{\geq T}$. Since there always exists $T\in\mathbb{I}_{\geq 1}$ such that $(\underline{\lambda}(\underline{G})-\overline{\lambda}(\overline{G})\eta^{T})>0$, Assumption 9 is satisfied, which finishes this proof. ∎
Proposition 10 essentially implies that state detectability (Assumptions 1) and uniform PE of the parameters (Assumption 9) is equivalent to uniform detectability (exponential i-IOSS) of the augmented state $x_{\mathrm{a}}^{\top}=[x^{\top},p^{\top}]$. Consequently, under these assumptions one could simply consider the augmented state $x_{\mathrm{a}}$ and apply MHE schemes and theory for state estimation (e.g., [2, 13, 27]). However, Assumption 9 is restrictive, usually not satisfied in practice, and its a priori verification is generally impossible. The proposed method from Section 3.1, on the other hand, provides a strict relaxation, since it is applicable in the practically relevant case where the parameters are only rarely (or never) excited (which violates Assumption 9 and hence implies that the augmented state cannot be uniformly detectable (i.e., exponentially i-IOSS) and no joint i-IOSS Lyapunov function satisfying (50) and (51) can exist).